OpenStudy (anonymous):

Which of the following equations has infinitely many real solutions? A. 5x + 1 = x - 1 B. 5(x - 5) = 5x + 25 C. 5(x - 5) = 5x - 25 D. 5x + 1 = 5x - 1

OpenStudy (anonymous):

@armyengineer225 im sorry for all of the questions i have only 6 more after this one, ur the one thats helped me most

OpenStudy (anonymous):

is it b? i dont know

OpenStudy (anonymous):

Infinitely many solutions means that 0 = 0. So on B, you have 5x-25 = 5x+25. 0 = 50, so that doesn't work. C comes out to 5x-25 = 5x-25, So 0 = 0. C is the answer.

OpenStudy (anonymous):

oh ok thank you i understand those kind of questions now!